two cars pas each other traveling in opposite directions on a highway, one is goig 1.5 times as fast as the other. at the end of 3.5 hours they are 192.5 miles apart. find each cars rate.

Answer by ankor@dixie-net.com(22740)

Two cars pass each other traveling in opposite directions on a highway, one is going 1.5 times as fast as the other. At the end of 3.5 hours, they are 192.5 miles apart. Find each cars rate.
:
Let s = speed of the slower car
Then
(1.5s) = speed of the faster car
:
Write a distance equation: Distance = time * speed.
:
They both will be traveling for 3.5 hr
;
Fast car dist + slow car dist = 192.5
3.5(1.5s) + 3.5s = 192.5
5.25s + 3.5s = 192.5
8.75s = 192.5
s = 192.5/8.75
s = 22 mph is the slow car
Then
1.5(22)) = 33 mph is the fast car
:
:
Check solution:
3.5(33) + 3.5(22) =
115.5 + 77 = 192.5: confirms out solution